OpenStudy (anonymous):

during ascent, and especially during decent, volume changes of trapped air in the middle ear can cause ear discomfort until the middle-ear pressure and exterior pressure are equalized. If the rapid descent rate of 7.0 m/s or faster commonly causes ear discomfort, what is the maximum rate of atmospheric pressure that is tolerable to most people?

OpenStudy (dboome):

Here's my reasoning, see what you think: We know that the pressure P at a height h in the atmosphere is given by\[P(h)=p_0e ^{-h/h_0}\] where: p = atmospheric pressure (measured in bars) h = height (altitude) p0 = is pressure at height h = 0 (surface pressure) h0 = scale height (which we know is about 7000m) Taking the derivative with respect to time of that, you get \[dP/dt = p_0*d(e ^{-h/h_0})/dt\]\[dP/dt = p_0\frac{ -\frac{ dh }{ dt} }{ h_0 }e ^{-h/h_0}\] However, with dh/dt given at -7m/s, h0 as 7km, h as 0 (is that reasonable?) and p0 = 1 bar, i arrive at the answer 1000bar/seconds, which doesn't seem plausible. Spot my errors(s)?

OpenStudy (anonymous):

Thank you so much! I do believe it requires a differential equation and that fits the bill.